I have the full atomic symbol 132Te2- and I'm having trouble figuring out the protons neutrons electrons for it. Also I have neutrons being 51 for a different problem, 37 electrons and a positive 3 charge but need help figuring out the protons for it and full atomic symbol.
1
Expert's answer
2017-05-01T09:59:08-0400
1. 132Te 2-

A = 132 – atomic mass

Z(p) = 52 protons (because of atomic number)

N(n) = A – Z = 132-52 = 80 Neutrons

E(e)= Z(p) = 52 electrons, but we have negative 2 charge. So we have 54

electrons.

Answer: 52 protons, 80 neutrons, 54 electrons.

2. N(n) = 51, E(e)=37, +3

Z(p) = E(e)+3 = 37+3=40 protons – atomic number

A = 40+51 = 91 – atomic mass

Zr

Answer: 40 protons, Zr.